# DHP: Discrete Hierarchical Planning for HRL Agents

### v2

@@ A.1 Policy Gradients for Trees
If $A(\tau)$ is any function independent of policy actions $z_{i}$, say $b(n_{i})$, then its net contribution to the policy gradient is 00.
@@
A.2 Policy Evaluation for Trees
We present the return and advantage estimation for trees as an extension of current return estimation methods for linear trajectories. As the return estimation for a state $s_{t}$ in linear trajectories depends upon the next state $s_{t+1}$, our tree return estimation method uses child nodes $(n_{2i+1},n_{2i+2})$ to compute the return for a node $n_{i}$. We extend the previous methods like lambda returns and Gen realized Advantage estimation (GAE) for trees.
The objective of our method is to reach nodes that are directly reachable. Such nodes are marked as terminal, and the agent receives a reward. For generalization, let’s say that when the agent takes an action $z_{i}$ at node $n_{i}$, it receives a pair rewards $(R_{2i+1}(n_{i},z_{i}),R_{2i+2}(n_{i},z_{i}))$ corresponding to the child nodes. Formally, the rewards $R(\tau)$ is an array of length equal to the length of the tree trajectory with $R_{0}=0$. Then, the agent’s task is to maximize the sum of rewards received in the tree trajectory $\mathbb{E}_{\tau}\sum_{i=0}^{\infty}R_{i}$. To consider future rewards, the returns for a trajectory can be computed as the sum of rewards discounted by their distance from the root node (depth), $\mathbb{E}_{\tau}\sum_{i=0}^{\infty}\gamma^{\lfloor\log_{2}(i+1)\rfloor-1}R_{i}$. Thus, the returns for each node can be written as the sum of rewards obtained and the discount-weighted returns thereafter:
@@ A.2 Policy Evaluation for Trees
Although this works theoretically, a flaw causes the agent to collapse to a degenerate local optimum. This can happen if the agent can generate a subgoal very similar to the initial or goal state ($\|s_{t},s_{\text{sub}}\|<\epsilon$ or $\|s_{g},s_{\text{sub}}\|<\epsilon$). A usual theme of reward systems for subgoal trees will be to have a high reward when the agent predicts a reachable or temporally close enough subgoal. Thus, if the agent predicts a degenerate subgoal, it receives a reward for one child node and the initial problem carries forward to the other node.
@@ A.2.1 Lambda Returns
Which essentially translates to the minimum of lambda returns using either of the child nodes as the next state. For theoretical generalization, note that the $\min$ operator in the return estimate is over the next states the agent is placed in. Thus, in the case of linear trajectories where there is only one next state, the $\min$ operator vanishes and the equation conveniently reduces to the standard return formulation for linear trajectories.
@@ Conclusion:
A similar argument can be shown for $\|\mathcal{T}^{1}V_{1}-\mathcal{T}^{1}V_{2}\|_{\infty}$ and $\|\mathcal{T}^{\lambda}V_{1}-\mathcal{T}^{\lambda}V_{2}\|_{\infty}$. Using the non-expansive property (Th. A.3) and absorbing the $\max$ operator with $\|\cdot\|_{\infty}$ leading to the standard form for linear trajectories.
@@
[…]
@@
$\frac{\gamma(1-\lambda)}{1-\gamma\lambda}<$ 11 11
@@
A.2.2 Bootstrapping with $D$-depth Returns
When the subtask tree branches end as terminal (or are masked as reachable), the agent receives a reward $=1$, which provides a learning signal using the discounted returns. However, when the branches do not end as terminal nodes, it does not provide a learning signal for the nodes above it, as the return is formulated as $\min$ of the returns from child nodes. In this case, we can replace the returns of the non-terminal leaf nodes with their value estimates. Therefore, in case the value estimate from the end node is high, indicating the agent knows how to solve the task from there onwards, it still provides a learning signal. The $n$-step return for a linear trajectory is written as:
@@ A.2.2 Bootstrapping with $D$-depth Returns
A.2.3 Properties of Tree Return Estimates
In section A.2 it can be seen how the tree return formulation for a node essentially reduces to the linear trajectory returns along the path of minimum return in the subtree under it. When the value function has reached the stationary point. For a subtask tree, if all branches end as terminal nodes, the return will be $\gamma^{D^{\prime}}1$, where $D^{\prime}$ is the depth of the deepest non-terminal node. Otherwise, it would be $\gamma^{D}V^{\prime}$ where $V^{\prime}$ is the non-terminal truncated node with the minimum return. Thus, it can be seen how higher depth penalizes the returns received at the root node with the discount factor $\gamma$. This property is true with the linear trajectories where the policy converges to the shortest paths to rewards to counter discounting [31, 28]. Thus, our goal-conditioned policy similarly converges to plans trees with minimum maximum depth: $\min_{\pi_{P}}(\max d_{i})$, where $d_{i}$ is the depth of node $n_{i}$.
This property also implies that the returns for a balanced tree will be higher than those for an unbalanced tree. The same sequence of leaf nodes can be created using different subtask trees. When the policy does not divide the task into roughly equally tough sub-tasks it results in an unbalanced tree. Since the tree is constrained to yield the same sequence of leaf nodes, its maximum depth $D_{U}$ will be higher than or equal to a balanced tree $D_{U}\geq D_{B}$. Thus, at optimality, the policy should subdivide the task in roughly equal chunks. However, it should be noted that two subtask trees with different numbers of leaf nodes can have the same maximum depth.
@@ B.1 Worker
B.2 Goal State Representations
Since the RSSM integrates a state representation using a sequence of observations, it does not work well for single observations. To generate goal state representations using single observations we train an MLP separately that tries to approximate the RSSM outputs ($s_{t}$) from the single observations ($o_{t}$). We call these representations static state representations. Moreover, since GCSR modules require state representations at large temporal distances, it can be practically infeasible to generate them using RSSM. Thus, we use static state representations to generate training data for the GCSR module as well. The MLP is a dense network with a $\tanh$ activation at the final output layer. It is trained to predict the RSSM output (computed using a sequence of images) using single image observations. To avoid saturating gradients, we use an MSE loss on the preactivation layer using labels transformed as $l_{\text{new}}=\text{atanh}(\text{clip}(l,\delta-1,1-\delta))$. The clipping helps avoid computational overflows, we use $\delta=10^{-4}$.
B.3 Implementation Details
We implement two functions: policy and train, using the hyperparameters shown in Table 2. The agent is implemented in Python/Tensorflow with XLA JIT compilation. Using XLA optimizations, the total training wall time is $2-3$ days on a consumer GPU (NVIDIA 4090 RTX 24gb). During inference, the compiled policy function runs in $4.6$ms on average, enabling real-time replanning at $217.39$ times per second (assuming minimal environmental overhead).
@@ B.3 Implementation Details
Worker abstraction length $K$ $8$
@@
Maximum Tree depth during training $D$ $5$
Maximum Tree depth during inference $D_{\text{Inf}}$ $8$
@@ B.3.1 Policy Function
Algorithm 1 DHP Policy Function
@@
3:     if mode == $eval$ then
@@
[…]
@@
18:  end if
19:  $a_{t}\sim\pi_{W}(a_{t}|s_{t},s_{wg})$
[…]
B.3.2 Train Function
The training function is executed every $16$-th step. A batch size $B$ of trajectories $\kappa$ and coarse trajectories $\kappa_{c}$ is sampled from the exploration trajectories or the expert data. The length of extracted trajectories is $L$ and the length of coarse trajectories is $L_{c}$ spanning over $L_{c}\times K$ time steps. Then the individual modules are trained sequentially (Alg. 6):
  • RSSM module is trained using $\kappa$ via the original optimization objective [13] followed by the static state representations (Sec. B.2).
@@ B.3.2 Train Function
Algorithm 2 Training the GCSR Module
0:  Experience dataset $data$
@@
[…]
@@
Algorithm 3 Training the Worker Policy
0:  Experience dataset $data$
[…]
2:  $\text{traj}\leftarrow\text{imagine}(\text{init},K)$On-policy trajectory rollout}
[…]
Algorithm 4 Training the Planner Policy
0:  Experience dataset $data$
@@
Algorithm 5 Training the Explorer Policy
0:  Experience dataset $data$
[…]
2:  $\text{traj}\leftarrow\text{imagine}(\text{init},H_{E})$On-policy trajectory rollout}
[…]
Algorithm 6 Overall Training Procedure
0:  Experience dataset $data$, current mode $mode$
@@ C.1 Positive Impacts
The imagination-based policy optimization mitigates hazards that can occur during learning. The efficient training can reduce the carbon footprint of the agents. The agent produces highly interpretable plans that can be verified before execution.
@@ Appendix D Sample Exploration Trajectories
We compare sample trajectories generated by various reward schemes in this section. Fig. 10 shows the sample trajectories of an agent trained to optimize the vanilla exploratory rewards. Fig. 11 shows the sample exploration trajectories of an agent that optimizes only for the GCSR-based rewards. Fig. 12 shows sample exploration trajectories of an agent that optimizes for GCSR exploratory rewards but without memory. The GCSR rewards-based agent generates trajectories that are less likely to lead to repeated path segments or remaining stationary. Removing the memory sometimes causes inefficient trajectories where the agent might not perform as expected.
Figure 10: Sample exploration trajectories using the vanilla exploratory rewards. The agent identifies states with unclear representations and successfully navigates to them, indicating sufficient navigation capabilities. However, once at the goal, the agent stays there, leading to lower-quality data.
@@ Appendix E Robo Yoga Samples
Figure 13: Figure shows a sample trajectory (every $8$-th frame) for the walker embodiment given a headstand goal. The agent can maintain a constant headstand; however, it sways about the goal position. This is because the agent is trained to reach the goal position and not to stay there.
@@
NeurIPS Paper Checklist
@@
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction make precise claims that align with the technical contributions and empirical validation:
    • Discrete rewards scheme performance (Fig. 8(a)).
    • Advantage estimation that generalizes beyond training depths with theoretical analysis (Fig. 8(b), Sec. A.2.3)
    • Exploration reward that encourages relevant trajectories (Fig. 8(c),11).
@@ 3. Answer: [Yes]
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: While we observe that our method succeeds in challenging visual goal conditioned planning scenarios, we have observed that it struggles in certain environments, like quadruped, and we have provided this result. Thus, our method may be suitable for certain environments, but not for all, and the related results have been provided.

### v3

@@ A.2 Policy Gradients for Trees
If $A(\tau)$ is any function independent of policy actions $z_{i}$, say $b(n_{i})$, then its net contribution to the policy gradient is 0.
@@
A.3 Policy Evaluation for Trees
We present the return and advantage estimation for trees as an extension of current return estimation methods for linear trajectories. As the return estimation for a state $s_{t}$ in linear trajectories depends upon the next state $s_{t+1}$, our tree return estimation method uses child nodes $(n_{2i+1},n_{2i+2})$ to compute the return for a node $n_{i}$. We extend the previous methods, like lambda returns and Gen realized Advantage estimation (GAE) for trees.
The objective of our method is to reach nodes that are directly reachable. Such nodes are marked as terminal, and the agent receives a reward. For generalization, let’s say that when the agent takes an action $z_{i}$ at node $n_{i}$, it receives a pair of rewards $(R_{2i+1}(n_{i},z_{i}),R_{2i+2}(n_{i},z_{i}))$ corresponding to the child nodes. Formally, the rewards $R(\tau)$ are an array of length equal to the length of the tree trajectory with $R_{0}=0$. Then, the agent’s task is to maximize the sum of rewards received in the tree trajectory $\mathbb{E}_{\tau}\sum_{i=0}^{\infty}R_{i}$. To consider future rewards, the returns for a trajectory can be computed as the sum of rewards discounted by their distance from the root node (depth), $\mathbb{E}_{\tau}\sum_{i=0}^{\infty}\gamma^{\lfloor\log_{2}(i+1)\rfloor-1}R_{i}$. Thus, the returns for each node can be written as the sum of rewards obtained and the discount-weighted returns thereafter:
@@ A.3 Policy Evaluation for Trees
Although this works theoretically, a flaw causes the agent to collapse to a degenerate local optimum. This can happen if the agent can generate a subgoal very similar to the initial or goal state ($\|s_{t},s_{\text{sub}}\|<\epsilon$ or $\|s_{g},s_{\text{sub}}\|<\epsilon$). A common theme in reward systems for subgoal trees is to have a high reward when the agent predicts a reachable or temporally close enough subgoal. Thus, if the agent predicts a degenerate subgoal, it receives a reward for one child node, and the initial problem carries forward to the other node.
@@ A.3.1 Lambda Returns
This essentially translates to the minimum of lambda returns using either of the child nodes as the next state. For theoretical generalization, note that the $\min$ operator in the return estimate is over the next states the agent is placed in. Thus, in the case of linear trajectories where there is only one next state, the $\min$ operator vanishes and the equation conveniently reduces to the standard return formulation for linear trajectories.
@@ Conclusion:
A similar argument can be shown for $\|\mathcal{T}^{1}V_{1}-\mathcal{T}^{1}V_{2}\|_{\infty}$ and $\|\mathcal{T}^{\lambda}V_{1}-\mathcal{T}^{\lambda}V_{2}\|_{\infty}$. Using the non-expansive property (Th. A.3) and absorbing the $\max$ operator with $\|\cdot\|_{\infty}$ leads to the standard form for linear trajectories.
@@
$\leq$ $\gamma(1-\lambda)\|V_{1}-V_{2}\|_{\infty}+\gamma\lambda\|\mathcal{T}^{\lambda}V_{1}-\mathcal{T}^{\lambda}V_{2}\|_{\infty}\quad\text{(Using triangle inequality)}$
@@
$\frac{\gamma(1-\lambda)}{1-\gamma\lambda}<$ $1$
@@
A.3.2 Bootstrapping with D-depth Returns
When the subtask tree branches end as terminal (or are masked as reachable), the agent receives a reward of $1$, which provides a learning signal using the discounted returns. However, when the branches do not end as terminal nodes, it does not provide a learning signal for the nodes above it, as the return is formulated as $\min$ of the returns from child nodes. In this case, we can replace the returns of the non-terminal leaf nodes with their value estimates. Therefore, in the case that the value estimate from the end node is high, indicating that the agent knows how to solve the task from that point onwards, it still provides a learning signal. The $n$-step return for a linear trajectory is written as:
@@ A.3.2 Bootstrapping with D-depth Returns
A.3.3 Properties of Tree Return Estimates
In section A.3 it can be seen how the tree return formulation for a node essentially reduces to the linear trajectory returns along the path of minimum return in the subtree under it. When the value function has reached the stationary point. For a subtask tree, if all branches end as terminal nodes, the return will be $\gamma^{D^{\prime}}1$, where $D^{\prime}$ is the depth of the deepest non-terminal node. Otherwise, it would be $\gamma^{D}V^{\prime}$ where $V^{\prime}$ is the non-terminal truncated node with the minimum return. Thus, it can be seen how higher depth penalizes the returns received at the root node with the discount factor $\gamma$. This property holds for linear trajectories, where the policy converges to the shortest paths to rewards, thereby counteracting discounting [sutton2018reinforcement, puterman2014markov]. Thus, our goal-conditioned policy similarly converges to plans trees with minimum maximum depth: $\min_{\pi_{P}}(\max d_{i})$, where $d_{i}$ is the depth of node $n_{i}$.
This property also implies that the returns for a balanced tree will be higher than those for an unbalanced tree. The same sequence of leaf nodes can be created using different subtask trees. When the policy does not divide the task into roughly equally tough sub-tasks it results in an unbalanced tree. Since the tree is constrained to yield the same sequence of leaf nodes, its maximum depth $D_{U}$ will be higher than or equal to a balanced tree $D_{U}\geq D_{B}$. Thus, at optimality, the policy should subdivide the task in roughly equal chunks. However, it is worth noting that two subtask trees with different numbers of leaf nodes can have the same maximum depth.
@@ B.1 Worker
B.2 Explorer
The ELBO objective for the unconditional state recall (UCSR) module is given as:
B.3 Goal State Representations
Since the RSSM integrates a state representation using a sequence of observations, it does not work well for single observations. To generate goal state representations using single observations, we train an MLP separately that tries to approximate the RSSM outputs ($s_{t}$) from the single observations ($o_{t}$). We refer to these representations as static state representations. Moreover, since GCSR modules require state representations at large temporal distances, it can be practically infeasible to generate them using RSSM. Thus, we use static state representations to generate training data for the GCSR module as well. The MLP is a dense network with a $\tanh$ activation at the final output layer. It is trained to predict the RSSM output (computed using a sequence of images) using single-image observations. To avoid saturating gradients, we use an MSE loss on the preactivation layer using labels transformed as $l_{\text{new}}=\text{atanh}(\text{clip}(l,\delta-1,1-\delta))$. The clipping helps avoid computational overflows; we use $\delta=10^{-4}$.
B.4 Implementation Details
We implement two functions: policy and train, using the hyperparameters shown in Table 3. The agent is implemented in Python/Tensorflow with XLA JIT compilation. Using XLA optimizations, the total training wall time is 232-3 days on a consumer GPU (NVIDIA 4090 RTX 24gb). During inference, the compiled policy function runs in 4.64.6ms on average, enabling real-time replanning at $217.39$ times per second (assuming minimal environmental overhead).
@@ B.4 Implementation Details
Worker abstraction length $K$ 88
@@
Maximum Tree depth during training $D$ 55
Maximum Tree depth during inference $D_{\text{Inf}}$ 88
@@ B.4.1 Policy Function
Algorithm 1 policy($o_{t}$, $o_{g}$, ($t$, $s_{t-1}$, $a_{t-1}$, $s_{wg}$, mem_buff, mode))
@@
3:  if mode == evaleval then
@@
13:   $\text{mem}_{t}\leftarrow\texttt{extract\_mem}(\text{mem\_buff})$ {Extract memory Sec. 2.4.2}
@@
17:  $\text{mem\_buff}\leftarrow\texttt{concat}(\text{mem\_buff}[1:],s_{t})$
18:end if
@@
20:return $a_{t}$, ($s_{t},s_{wg},\text{mem\_buff}$)
B.4.2 Train Function
The training function is executed every $16$-th step. A batch size $B$ of trajectories $\kappa$ and coarse trajectories $\kappa_{c}$ is sampled from the exploration trajectories or the expert data. The length of extracted trajectories is $L$ and the length of coarse trajectories is $L_{c}$ spanning over $L_{c}\times K$ time steps. Then the individual modules are trained sequentially, shown in the overall script as Alg. 6:
  • RSSM module is trained using $\kappa$ via the original optimization objective [hafner2019learning] followed by the static state representations (Sec. B.3).
@@ B.4.2 Train Function
Algorithm 2 train_gcsr(datadata)
0: Experience dataset datadata
@@
3:  Append $\text{extract\_triplets}(data,q)$ to triplets {Extract $(s_{t},s_{t+q/2},s_{t+q})$}
@@
Algorithm 3 train_worker(datadata)
0: Experience dataset datadata
@@
Algorithm 4 train_planner(datadata)
0: Experience dataset datadata
@@
Algorithm 5 train_explorer
0: Experience dataset datadata
[…]
Algorithm 6 Overall Algorithm
0: Environment envenv, DHP Agent agentagent, Replay buffer bufferbuffer
[…]
4:while $t<5\times 10^{6}$ do
5:  if $t<3\times 10^{6}$ then
6:   mode = evaleval
7:  else
8:   mode = exploreexplore
9:  end if
10:  $a_{t},(s_{t},s_{wg}$, mem_buff) $\leftarrow agent.\texttt{policy}(o_{t},o_{g},(t,s_{t-1},a_{t-1},s_{wg}$, mem_buff, mode)) {Act using Alg. 1}
11:  $o_{t+1},r_{t}\leftarrow env.\texttt{step}(a_{t})$
12:  $buffer.\texttt{store}(o_{t},a_{t},r_{t})$ {Store step in the replay buffer}
13:  if $t\mod 16==0$ then
14:   $data\leftarrow buffer.\texttt{sample\_batch}()$ {Sample a batch of trajectories from the buffer}
15:   $agent.\texttt{train\_rssm}(data)$ {Train world model per [hafner2019learning]}
16:   $agent.\texttt{train\_gcsr}(data)$ {Alg. 2}
17:   $agent.\texttt{train\_worker}(data)$ {Alg. 3}
18:   $agent.\texttt{train\_planner}(data)$ {Alg. 4}
19:   $agent.\texttt{train\_explorer}(data)$ {Alg. 5}
20:  end if
21:  $s_{t-1},a_{t-1},t\leftarrow s_{t},a_{t},t+1$ {Update recurrent variables}
22:end while
@@ C.1 Positive Impacts
The imagination-based policy optimization mitigates hazards that can occur during learning. Efficient training can reduce the carbon footprint of the agents. The agent produces highly interpretable plans that can be verified before execution.
@@ Appendix D Sample Exploration Trajectories
We compare sample trajectories generated by various reward schemes in this section. Fig. 10 shows the sample trajectories of an agent trained to optimize the vanilla exploratory rewards. Fig. 11 shows the sample exploration trajectories of an agent that optimizes only for the GCSR-based rewards. Fig. 12 shows sample exploration trajectories of an agent that optimizes for GCSR exploratory rewards but without memory. The GCSR rewards-based agent generates trajectories that are less likely to lead to repeated path segments or remaining stationary. Removing the memory can sometimes cause inefficient trajectories, where the agent may not perform as expected.
Figure 10: Sample exploration trajectories using the vanilla exploratory rewards. The agent identifies states with unclear representations and successfully navigates to them, indicating sufficient navigation capabilities. However, once the agent reaches the goal, it stays there, leading to lower-quality data.
@@ Appendix E Robo Yoga Samples
Figure 13: Figure shows a sample trajectory (every 88-th frame) for the walker embodiment given a headstand goal. The agent can maintain a constant headstand; however, it sways about the goal position. This is because the agent is trained to reach the goal position and not to stay there.
@@
Appendix F Offline DHP
While the online formulation presented in earlier sections is highly flexible and requires minimal external dependencies, scaling to extremely long-horizon and complex tasks (such as OGBench [park2024ogbench]) reveals practical challenges:
  • Compounding errors in subgoal prediction: Since subgoals at lower tree levels depend on predictions at higher levels, any hallucinations in the CVAE-based subgoal generation accumulate with depth (Figure 14 shows examples). This becomes problematic when planning over very long horizons where recursive errors compound.
  • Long-horizon exploration inefficiency: Undirected exploration can require impractically long training times in large environments. Moreover, many states that yield high exploratory rewards may be irrelevant to the external task (e.g., an agent lying upside-down throughout a maze provides novel states but unhelpful coverage).
To address these limitations while preserving our core contributions—discrete reachability checks and tree-structured return estimation—we develop a simplified offline variant of DHP. Demonstrating that our planning principles are architecture-agnostic and applicable across different settings.
F.1 Architecture
The offline agent maintains our core planning objectives while replacing components that proved problematic in complex domains:
@@ F.1 Architecture
    Goal representation network $\phi(s_{t},s_{g})\rightarrow z$ (similar to HIQL [park2023hiql]): Replaces world-model-based state representations with a learned encoder that directly maps state-goal pairs to a length-normalized representation space.
@@
    Hierarchical policies:
    • Manager $\pi_{M}(s_{t},s_{g})\rightarrow z$: Predicts optimal subgoal representations in the learned $z$-space.
    • Worker $\pi_{W}(s_{t},z)\rightarrow a_{t}$: Executes low-level actions to reach subgoals.